# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v2

@@ 3 Theoretical Methods to Obtain Gravitational Lensing Observables
This section gives descriptions on the theoretical methods used to calculate the gravitational lensing observables in our work. The treatments of particle orbits (which are geodesics) are presented in subsection 3.1. Schemes to obtained gravitational deflection angle of light and time delay of light for finite distance light source and observer are introduced in subsections 3.2 - 3.3. Finally, the formulations on the trajectories and precession angle for massive object’s bound orbits around black holes are given in subsection 3.4.
@@ 3.1 Geodesics
In the gravitational filed, massive and massless particles are moving along geodesics in 4 dimensional spacetime. Here we provide a derivation of the reduced geodesic equation, which turns out to be extremely useful in the exploration of gravitational lensing observables. For a spherically symmetric spacetime with the metric form
$d\tau^{2}=f(r)dt^{2}-\frac{1}{f(r)}dr^{2}-r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$ (8)
the following conserved quantities can be introduced from the symmetry and Killing vectors of spacetime
$J$$\equiv$ $r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$ (9a)
$E$$\equiv$ $f(r)\frac{dt}{d\lambda}$ (9b)
$\epsilon$$\equiv$ $g_{\mu\nu}dx^{\mu}dx^{\nu}=f(r)\bigg{(}\frac{dt}{d\lambda}\bigg{)}^{2}-\frac{1}{f(r)}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}-r^{2}\bigg{(}\frac{d\theta}{d\lambda}\bigg{)}^{2}-r^{2}\sin^{2}\theta\bigg{(}\frac{d\phi}{d\lambda}\bigg{)}^{2}$ (9c)
Here, $\lambda$ is an affine parameter in a particle orbit, $J$ is the conserved angular momentum along the particle orbit, and $E^{2}/2$ can be viewed as the conserved energy along a particle orbit. Specifically, for spherically symmetric spacetime, we can always assume that test particles are moving in the equatorial plane $\theta=\pi/2$, and the reduced differential equation can be obtained using these conserved quantities
$\frac{1}{2}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+\frac{1}{2}f(r)\bigg{[}\frac{J^{2}}{r^{2}}+\epsilon\bigg{]}=\frac{1}{2}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+V_{\text{eff}}(r)=\frac{1}{2}E^{2}$ (10)
@@ 3.2 Gravitational Deflection Angle of Light
The gravitational deflection angle reflects the bending or distortion effects on particle orbit in a gravitational field. In the gravitational lensing observation, when the light source and observer are located at a finite distance region to the central black hole (with the radial coordinate $r=r_{\text{S}}$ and $r=r_{\text{O}}$, respectively), the gravitational deflection angle of light is defined through [106]
$\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$ (11)
@@
$\sin\Psi=\frac{b\cdot\sqrt{f(r)}}{r}$ (12)
The definition of gravitational deflection angle of light in expression (11) for finite distance light source and observer can be understood from an instructive way using thin lens approximation. In the thin lens approximation, the gravitational deflection angle for finite distance light source and observer is illustrated in figure 1. The thin lens approximation assumes that the spacetime is flat everywhere except for the location of gravitational thin lens (the central black hole position). For the quadrangle depicted in figure 1, the exterior angles at each point are $\Psi_{\text{S}}$, $\alpha$, $\pi-\Psi_{\text{O}}$, $\pi-\Delta\phi_{\text{OS}}$, respectively. Under the thin lens approximation, the sum of exterior angles for the quadrangle equals $2\pi$, which justify the definition of gravitational deflection angle $\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$. However, it should be noted that this definition in expression (11) given by Ishihara et al. is not restricted in the thin lens approximation. It is a consistent and proper definition of gravitational deflection angle that are hold in more generally gravitational lensing treatments (which can go beyond the thin lens approximation), see references [106, 107, 108, 109, 110] for more discussions on this issue.
@@
$\frac{d\phi}{dr}=\frac{d\phi}{d\lambda}\cdot\frac{d\lambda}{dr}=\pm\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}$ (13)
where we have used the affine parameter $\epsilon=0$ for massless photon, the conserved energy parameter $E=f(r)\frac{dt}{d\lambda}$, conserved angular momentum $J=r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$, and the definition of impact parameter $b\equiv|J/E|$. The plus and minus sign $\pm$ can be determined in the following way. When the photon beams travel along a scattering orbit from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as the azimuthal angle $\phi$ increases, until photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase with the increasing of azimuthal angle. This we derive the following relation
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}<0\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point$r=r_{0}$}$
$\frac{d\phi}{dr}=\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}>0\ \ \ \ \ \ \ \text{from tuning point $r=r_{0}$ to observer position$r=r_{\text{O}}$}$
@@
$\Rightarrow$ $b^{2}=\frac{J^{2}}{E^{2}}=\frac{r_{0}^{2}}{f(r_{0})}$
@@
With the expression of angle $\Psi$ in equation (12) and the variation of azimuthal angle, the gravitational deflection angle of light can be expressed as
@@
$=$ $\arcsin\bigg{(}\frac{b\sqrt{f(r_{\text{R}})}}{r_{\text{R}}}\bigg{)}+\arcsin\bigg{(}\frac{b\sqrt{f(r_{\text{S}})}}{r_{\text{S}}}\bigg{)}-\pi+\int_{r_{0}}^{r_{\text{S}}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}$
@@ 3.3 Time Delay of Light
In the gravitational field, the elapse of time (measured by an observer) during the light propagation can be strongly affected by gravitation, and this is the time delay effect of gravitational field. To calculate the time delay of light in gravitational lensing observations, it is necessary to compute the variation of time coordinate $t$ (measured by an distant observer) in the photon orbit. For the massless photon traveling along null geodesics (with $\epsilon=0$), one can derive the time derivative of radial coordinate from the reduced geodesic equation in (10)
$\frac{dr}{dt}=\frac{dr}{d\lambda}\cdot\frac{d\lambda}{dt}=\pm f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}$ (17)
The plus and minus sign $\pm$ can be determined similar to the calculation of gravitational deflection angle. When the photon beams are emitted from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as time passes, until the photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase as time advances. So we have the relation
$\frac{dt}{dr}=-\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}<0\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}$
$\frac{dt}{dr}=\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}>0\ \ \ \ \ \ \ \text{from tuning point $r=r_{0}$ to observer position $r=r_{\text{O}}$}$
In the gravitational lensing observation, when the light source and observer are located at $r=r_{\text{S}}$ and $r=r_{\text{O}}$, the time delay of light during the light propagation process can be expressed as
@@
$=$ $\int_{r_{0}}^{r_{\text{S}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}-\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}-\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$
@@
Furthermore, it should be addressed that the time delay of light calculated using expression (19) is the time delay measured by the conventional static time coordinate $t$ in spacetime metric expression (8). For an asymptotically flat spacetime, the static time coordinate corresponds to the time measured by a very distant observer (at the asymptotically flat region). On the other hand, if one wants to calculate the time delay measured by the proper time of an finite distance static observer, a redshift factor should be taken into account
$\Delta T_{\text{static observer's proper time unit}}=\frac{\Delta T}{\text{redshift factor}}=\frac{\Delta T}{\sqrt{f(r_{\text{O}})}}$ (20)
with $r_{\text{O}}$ to be the position of finite distance static observer.
@@ 3.4 Massive Object’s Orbits and Precession Angle
Besides the gravitational deflection angle and time delay of light discussed in subsections 3.2-3.3 (which is relevant to a scattering orbit for massless photons), the massive object’s orbits are also greatly influenced by gravitational field, both for bound orbits and scattering orbits. Notably, massive object’s bound orbits may exhibit precession. In the gravitational lensing observations, the precession of bound orbits is frequently used to test the gravitational field around the supermassive black hole in galaxy center (such as the observations relevant to Sgr*A in Milky Way Galaxy).
@@
$\bigg{(}\frac{dr}{d\tau}\bigg{)}^{2}$ $=$ $E^{2}-V_{\text{eff}}(r)=E^{2}-f(r)\bigg{[}\frac{J^{2}}{r^{2}}+1\bigg{]}$ (21a)
$\frac{d\phi}{d\tau}$ $=$ $\frac{J}{r^{2}}$ (21b)
@@
$\Rightarrow$ $\bigg{(}\frac{du}{d\phi}\bigg{)}^{2}=\frac{1}{b^{2}}-F(u)\bigg{[}u^{2}+\frac{1}{J^{2}}\bigg{]}$
where we have used the new variable $u=1/r$ in the geodesic equation. Differentiating each side of this equation with respect to $\phi$, a second order differential equation for massive object’s orbit can be obtained
$\frac{d^{2}u}{d\phi^{2}}=-\frac{1}{2}\cdot\frac{dF(u)}{du}\cdot\bigg{[}u^{2}+\frac{1}{J^{2}}\bigg{]}-F(u)\cdot u$ (23)
The function $F(u)=f(1/r)$ is the spacetime metric component expressed using new variable $u=1/r$. Eventually, the massive object’s orbit is determined by numerically solving the differential equation (23).
@@
An important feature of the massive object’s bound object is the precession of azimuthal angle $\phi$, which is illustrated in figure 2. Given a bound timelike geodesic orbit for a massive object, there are maximal and minimal distance positions to central black hole in this orbit, denoting as $r_{\text{max}}$ and $r_{\text{min}}$. The precession of azimuthal angle $\phi$ in an orbit period can be calculated as
@@
$=$ $2\int_{r_{\text{min}}}^{r_{max}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}-2\pi$
In the derivation of precession angle $\Delta\phi_{\text{precession}}$, we have used the derivative of of azimuthal angle with respect to radial coordinate for a timelike geodesic orbit
$\frac{d\phi}{dr}=\frac{d\phi}{d\tau}\cdot\frac{d\tau}{dr}=\pm\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}$ (25)
The plus and minus sign $\pm$ can be determined as follows. When a massive object begins to move along timelike geodesics from the maximal distance position $r_{\text{max}}$, the radial coordinate $r$ decreases as azimuthal angle $\phi$ increases, until the massive object reaches the minimal distance position $r=r_{\text{min}}$ to central supermassive black hole. After passing the minimal distance position $r=r_{\text{min}}$, the massive object’s radial coordinate starts to increase as azimuthal angle increases. Thus we can get the relation
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}<0\ \ \ \text{from maximal distance $r=r_{\text{max}}$ to minimal distance $r=r_{\text{min}}$}$
$\frac{d\phi}{dr}=\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}>0\ \ \ \ \ \text{from minimal distance $r=r_{\text{min}}$ to maximal distance $r=r_{\text{max}}$}$
Furthermore, the maximal and minimal distance position to the central black hole in a massive object’s bound orbit can be determined via
$\frac{dr}{d\phi}\bigg{|}_{r=r_{\text{max}},\ r=r_{\text{min}}}=r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}\bigg{|}_{r=r_{\text{max}},\ r=r_{\text{min}}}=0$ (27)

### v3

@@ 3 Theoretical Method
This section gives descriptions on the theoretical methods used to calculate the gravitational lensing observables, massive object’s bound orbit, and black hole shadow. The treatments of particle orbits (which are geodesics) are presented in subsection 3.1. Schemes to obtain gravitational deflection angle of light and time delay of light for finite distance light source and observer are introduced in subsections 3.2 - 3.3. The method on the trajectories and precession angle for massive object’s bound orbits around black holes is reviewed in subsection 3.4. Finally, the approach to obtain black hole shadow is given in subsection 3.5.
@@ 3.1 Geodesics
In the gravitational field, massive and massless particles are moving along geodesics in 4 dimensional spacetime. Here we provide a derivation of the reduced geodesic equation, which turns out to be extremely useful in the exploration of gravitational lensing observables. For a spherically symmetric spacetime with the metric form
$d\tau^{2}=f(r)dt^{2}-\frac{1}{f(r)}dr^{2}-r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (8)
the following conserved quantities can be introduced from the symmetry and Killing vectors of spacetime [136, 137]
$J$$\equiv$ $r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda},$ (9a)
$E$$\equiv$ $f(r)\frac{dt}{d\lambda},$ (9b)
$\epsilon$$\equiv$ $g_{\mu\nu}dx^{\mu}dx^{\nu}=f(r)\bigg{(}\frac{dt}{d\lambda}\bigg{)}^{2}-\frac{1}{f(r)}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}-r^{2}\bigg{(}\frac{d\theta}{d\lambda}\bigg{)}^{2}-r^{2}\sin^{2}\theta\bigg{(}\frac{d\phi}{d\lambda}\bigg{)}^{2}.$ (9c)
Here, $\lambda$ is an affine parameter in a particle orbit, $J$ is the conserved angular momentum along the particle orbit, and $E^{2}/2$ can be viewed as the conserved energy along a particle orbit. Specifically, for spherically symmetric spacetime, we can always assume that test particles are moving in the equatorial plane $\theta=\pi/2$, and the reduced differential equation can be obtained using these conserved quantities [136, 137]
$\frac{1}{2}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+\frac{1}{2}f(r)\bigg{[}\frac{J^{2}}{r^{2}}+\epsilon\bigg{]}=\frac{1}{2}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+V_{\text{eff}}(r)=\frac{1}{2}E^{2}.$ (10)
@@ 3.2 Gravitational Deflection Angle of Light
In astrophysical observations, photons beams emitted from distant luminous sources can be deflected and distorted due to the gravitational field generated by extremely massive celestial bodies (such as supermassive black holes at the center of galaxies, or other huge massive gravitational sources). The gravitational deflection angle, which quantifies the change of tangent direction during the photon propagation, directly reflects the distortion effects induced by a gravitational field on photon (or other particle) trajectories. The larger deflection angle of light corresponds to stronger deflection and distortion effects of gravitation during the photon propagation process.
In the gravitational lensing observation, when light source and observer are located at a finite distance region to the central black hole (with the radial coordinate $r=r_{\text{S}}$ and $r=r_{\text{O}}$, respectively), the gravitational deflection angle of light is defined through [139]
$\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}.$ (11)
@@
$\sin\Psi=\frac{b\cdot\sqrt{f(r)}}{r}.$ (12)
The definition of gravitational deflection angle of light in expression (11) for finite distance light source and observer can be understood from an instructive way using thin lens approximation. In the thin lens approximation, the gravitational deflection angle for finite distance light source and observer is illustrated in figure 1. The thin lens approximation assumes that the spacetime is flat everywhere except for the location of gravitational thin lens (the central black hole position). For the quadrangle depicted in figure 1, the exterior angles at each point are $\Psi_{\text{S}}$, $\alpha$, $\pi-\Psi_{\text{O}}$, $\pi-\Delta\phi_{\text{OS}}$, respectively. Under the thin lens approximation, the sum of exterior angles for the quadrangle equals $2\pi$, which justify the definition of gravitational deflection angle $\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$. However, it should be noted that this definition in expression (11) given by Ishihara et al. is not restricted in the thin lens approximation. It is a consistent and proper definition of gravitational deflection angle, which is hold in more general gravitational lensing treatments (it can go beyond the thin lens approximation), see references [139, 140, 141, 142, 143] for more discussions on this issue.
@@
$\frac{d\phi}{dr}=\frac{d\phi}{d\lambda}\cdot\frac{d\lambda}{dr}=\pm\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}},$ (13)
where we have used the affine parameter $\epsilon=0$ for massless photon, the conserved energy parameter $E=f(r)\frac{dt}{d\lambda}$, conserved angular momentum $J=r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$, and the definition of impact parameter $b\equiv|J/E|$. The plus and minus signs $\pm$ can be determined in the following way. When the photon beams travel along a scattering orbit from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as the azimuthal angle $\phi$ increases, until photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase with the increasing of azimuthal angle. Thus, we derive the following relation
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}<0,\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}$
$\frac{d\phi}{dr}=\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}>0.\ \ \ \ \ \ \ \text{from tuning point $r=r_{0}$ to observer position $r=r_{\text{O}}$}$
@@
$\Rightarrow$ $b^{2}=\frac{J^{2}}{E^{2}}=\frac{r_{0}^{2}}{f(r_{0})}.$
@@
With the expression of angle $\Psi$ in equation (12) and the variation of azimuthal angle, the gravitational deflection angle of light can be expressed as [139, 137]
@@
$=$ $\arcsin\bigg{(}\frac{b\sqrt{f(r_{\text{R}})}}{r_{\text{R}}}\bigg{)}+\arcsin\bigg{(}\frac{b\sqrt{f(r_{\text{S}})}}{r_{\text{S}}}\bigg{)}-\pi+\int_{r_{0}}^{r_{\text{S}}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}.$
@@ 3.3 Time Delay of Light
In the gravitational field, the elapse of time (measured by an observer) during the light propagation can be strongly affected by gravitation, and this is the time delay effect of gravitational field. To calculate the time delay of light in gravitational lensing observations, it is necessary to compute the variation of time coordinate $t$ (measured by a distant observer) in the photon orbit. For the massless photon traveling along null geodesics (with $\epsilon=0$), one can derive the time derivative of the radial coordinate from the reduced geodesic equation in (10)
$\frac{dr}{dt}=\frac{dr}{d\lambda}\cdot\frac{d\lambda}{dt}=\pm f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}.$ (17)
The plus and minus signs $\pm$ can be determined similar to the calculation of gravitational deflection angle. When the photon beams are emitted from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as time passes, until the photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase as time advances. So we have the relation
$\frac{dt}{dr}=-\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}<0,\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}$
$\frac{dt}{dr}=\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}>0.\ \ \ \ \ \ \ \text{from tuning point $r=r_{0}$ to observer position $r=r_{\text{O}}$}$
In the gravitational lensing observation, when the light source and observer are located at $r=r_{\text{S}}$ and $r=r_{\text{O}}$, the time delay of light during the light propagation process can be expressed as [137, 138]
@@
$=$ $\int_{r_{0}}^{r_{\text{S}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}-\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}-\sqrt{r_{\text{O}}^{2}-r_{0}^{2}},$
@@
Furthermore, it should be addressed that the time delay of light calculated using expression (19) is the time delay measured by the conventional static time coordinate $t$ in spacetime metric expression (8). For an asymptotically flat spacetime, the static time coordinate corresponds to the time measured by a very distant observer (at the asymptotically flat region). On the other hand, if one wants to calculate the time delay measured by the proper time of a finite distance static observer, a redshift factor should be taken into account
$\Delta T_{\text{static observer's proper time unit}}=\frac{\Delta T}{\text{redshift factor}}=\frac{\Delta T}{\sqrt{f(r_{\text{O}})}},$ (20)
with $r_{\text{O}}$ to be the position of the finite distance static observer.
@@ 3.4 Massive Object’s Orbits and Precession Angle
Besides the gravitational deflection angle and time delay of light discussed in subsections 3.2-3.3 (which is relevant to a scattering orbit for massless photons), the massive object’s orbits are also greatly influenced by the gravitational field, both for bound orbits and scattering orbits. Notably, massive object’s bound orbits may exhibit precession. In astrophysical observations, the precession of bound orbits is frequently used to test the gravitational field around the supermassive black hole in a galaxy center (such as the observations relevant to Sgr*A in Milky Way Galaxy).
@@
$\bigg{(}\frac{dr}{d\tau}\bigg{)}^{2}$ $=$ $E^{2}-V_{\text{eff}}(r)=E^{2}-f(r)\bigg{[}\frac{J^{2}}{r^{2}}+1\bigg{]},$ (21a)
$\frac{d\phi}{d\tau}$ $=$ $\frac{J}{r^{2}},$ (21b)
@@
$\Rightarrow$ $\bigg{(}\frac{du}{d\phi}\bigg{)}^{2}=\frac{1}{b^{2}}-F(u)\bigg{[}u^{2}+\frac{1}{J^{2}}\bigg{]},$
where we have used a new variable $u=1/r$ in the geodesic equation. Differentiating each side of this equation with respect to $\phi$, a second-order differential equation for massive object’s orbit can be obtained [136, 137]
$\frac{d^{2}u}{d\phi^{2}}=-\frac{1}{2}\cdot\frac{dF(u)}{du}\cdot\bigg{[}u^{2}+\frac{1}{J^{2}}\bigg{]}-F(u)\cdot u\ .$ (23)
The function $F(u)=f(1/r)$ is the spacetime metric component expressed using the new variable $u=1/r$. Eventually, the massive object’s orbit is determined by numerically solving the differential equation (23).
@@
An important feature of the massive object’s bound object is the precession of azimuthal angle $\phi$, which is illustrated in figure 2. Given a bound timelike geodesic orbit for a massive object, there are maximal and minimal distance positions to central black hole in this orbit, denoting as $r_{\text{max}}$ and $r_{\text{min}}$. The precession of azimuthal angle $\phi$ in an orbit period can be calculated as [137, 138]
@@
$=$ $2\int_{r_{\text{min}}}^{r_{max}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}-2\pi.$
In the derivation of precession angle $\Delta\phi_{\text{precession}}$, we have used the derivative of azimuthal angle with respect to radial coordinate for a timelike geodesic orbit
$\frac{d\phi}{dr}=\frac{d\phi}{d\tau}\cdot\frac{d\tau}{dr}=\pm\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}.$ (25)
The plus and minus signs $\pm$ can be determined as follows. When a massive object begins to move along timelike geodesics from the maximal distance position $r_{\text{max}}$, the radial coordinate $r$ decreases as the azimuthal angle $\phi$ increases, until the massive object reaches the minimal distance position $r=r_{\text{min}}$ to central supermassive black hole. After passing the minimal distance position $r=r_{\text{min}}$, the massive object’s radial coordinate starts to increase as the azimuthal angle increases. Thus, we can get the relation
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}<0,\ \ \ \text{from maximal distance $r=r_{\text{max}}$ to minimal distance $r=r_{\text{min}}$}$
$\frac{d\phi}{dr}=\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}>0.\ \ \ \ \ \text{from minimal distance $r=r_{\text{min}}$ to maximal distance $r=r_{\text{max}}$}$
Furthermore, the maximal and minimal distance positions to the central black hole in a massive object’s bound orbit can be determined via
$\frac{dr}{d\phi}\bigg{|}_{r=r_{\text{max}},\ r=r_{\text{min}}}=r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}\bigg{|}_{r=r_{\text{max}},\ r=r_{\text{min}}}=0.$ (27)
3.5 Black Hole Shadow
This subsection presents the theoretical method to obtain the black hole shadow size. The critical shadow radius of a black hole is determined by the unstable photon sphere, which can be calculated from the effective potential of photons moving in a gravitational field, or through a geometric approach proposed in a recent works [144]. In this work, we calculate the position of unstable photon sphere and the black hole shadow radius based on the geometric approach, in which the photon sphere and its stability are uniquely determined from Gaussian curvature and geodesic curvature in the optical geometry of black hole spacetime. Firstly, any photon orbits maintain their geodesic nature when they are transformed into optical geometry, so the geodesic curvature for a photon sphere vanishes. This geodesic curvature condition for photon sphere is completely equivalent to find the local extreme point of effective potential [144, 145, 146]
$\kappa_{g}(r=r_{\text{ph}})=0\ \ \Leftrightarrow\ \ \frac{dV_{\text{eff}}(r)}{dr}\bigg{|}_{r=r_{\text{ph}}}=0.$ (28)
$\mathcal{K}(r=r_{\text{unstable}})<0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{unstable}}}<0,$ (29a)
$\mathcal{K}(r=r_{\text{stable}})>0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{stable}}}>0.$ (29b)
The absolute apparent size of black hole shadow viewed by an observer located at infinity (the asymptotically flat region) is characterized by the critical shadow radius. With the unstable photon sphere obtained in equations (28) and (29a), the critical shadow radius of a black hole can be calculated through the relation
$r_{sh}=b_{\text{critical}}=\frac{r_{\text{unstable}}}{\sqrt{f(r=r_{\text{unstable}})}}.$ (30)